OpenStudy (lgbasallote):

is this rational? \[\huge \frac{1.5}{2.0}\]

OpenStudy (anonymous):

Yes...1.5/2.0 = 3/4 isn't that rational?

OpenStudy (lgbasallote):

wait...let me think of a different fraction that doesn't become integers...

OpenStudy (anonymous):

I mean the represnted form is not rational but the simplified form is...

OpenStudy (lgbasallote):

\[\huge \frac{1.52}{1.63}\] i doubt that can be turned to integers

OpenStudy (lgbasallote):

so is that rational?

OpenStudy (anonymous):

Multiply by 100

OpenStudy (anonymous):

p/q where p and q are integers and q is not equal to zero

OpenStudy (anonymous):

And divide by 100

OpenStudy (anonymous):

152/163

OpenStudy (lgbasallote):

hmm good point...

OpenStudy (anonymous):

Again..the given form is not rational..

OpenStudy (lgbasallote):

\[\huge \frac{\sqrt 2}{\sqrt 3}\] ////

OpenStudy (lgbasallote):

rational?

OpenStudy (anonymous):

1.5/2 is not a rational number but 3/4 is...

OpenStudy (lgbasallote):

1.5/2 is actually rational

OpenStudy (anonymous):

No..1.5 is not an integer...

OpenStudy (anonymous):

root2 by root3 is not rational.

OpenStudy (anonymous):

p/q will be called a rational number if p and q are integers only.

OpenStudy (lgbasallote):

a rational number is any number that *can* be expressed as quotient of two integers...note.. *can*

OpenStudy (lgbasallote):

that was why i was trying to think of a number that cannot be simplified to two integers but still rational

OpenStudy (anonymous):

Okay got it...

OpenStudy (anonymous):

I never really stressed on "can"....I don't know if I really knew that atleast during my primary school!!!

OpenStudy (lgbasallote):

anyway.... so you tell me @sauravshakya that the quotient of two irrational numbers don't equa; to rational?

OpenStudy (anonymous):

 is rational